OpenStudy (anonymous):

Find the y coordinates for (-8, ), (-4, ), (4, ) y=3/4x-3

OpenStudy (anonymous):

plug in the known values of x that are given into the formula. y=(3/4)(-8)-3=-9 for the first one.

OpenStudy (anonymous):

y = (3/4)(8)-3 = 3 y = (3/4)(-4)-3 = -6 y = (3/4)(4)-3 = 0

OpenStudy (anonymous):

easy way in factoring?